# Cluster-cyclic condition of skew-symmetrizable matrices of rank 3 via the Markov constant

### v1

@@ 3. Markov constant
Suppose that $p_{n}$ converges. This is equivalent that $f_{2k+1}$ ($k\in\mathbb{Z}_{>0}$) converges. Let $n=2k+1$. When $r=2$, it holds that $\theta=0$. Since $u_{n}(2)=U_{n}(1)=n+1$, we have $f_{n}(S)=q(n+1)-pn=(q-p)n+q$. Thus, if $f_{n}(S)$ converges, we have $q-p=qe^{0}-p=0$. By considering the maps $\gamma_{1}$ and $\gamma_{2}$, we may obtain $p_{n}=q_{n}=p=q$ for any $n\in\mathbb{Z}_{>0}$. If $r>2$, or equivalently, if $\theta>0$, it holds that $\sin n\theta>0$ for any $n\in\mathbb{Z}_{>0}$ and we have $\frac{\sinh(n+1)\theta}{\sinh n\theta}=\frac{e^{(n+1)\theta}-e^{-(n+1)\theta}}{e^{n\theta}-e^{-n\theta}}=\frac{e^{\theta}-e^{-(2n+1)\theta}}{1-e^{-2n\theta}}\to e^{\theta}$ when $n\to\infty$. Since $u_{n}(r)=U_{n}(\cosh\theta)=\frac{\sinh(n+1)\theta}{\sinh n\theta}$, we have
@@
Unless $qe^{\theta}-p\neq 0$, it does not converge since $\sinh(2n+1)\theta\to\infty$ when $n\to\infty$. When $p=qe^{\theta}$, we have
@@
Later, we will show $C(S)\leq 4$ for any cluster-positive element $S$. It was shown in [FT19] by using the geometrical aspects of mutations. Here, we give a proof which is more similar to that in [BBH11], althoug we need some modifications for the real case.
@@
(3) A triplet $S$ satisfies (M3) if and only if $2q>pr$. In this case, $S$ satisfies $S\nleq\gamma_{1}(S),\gamma_{2}(S)$.
@@
Without loss of generality, we may assume $p\geq q\geq r>0$. By Proposition 7.3 (3), we have $2q>pr$ and $2p>qr$. By multiplying the corresponding sides, we have
@@
(a) We have three inequarities $p,q,r\geq 2$.
@@
which satisfies $t_{i-1}\neq t_{i}$ ($i=1,2,\dots,k$), then we have $S_{i-1}\leq S_{i}$. In particular, for any $S^{\prime}\in\Gamma(S)$, there is a sequence (46) with $S_{i-1}\leq S_{i}$, and $S_{i-1}\neq S_{i}$ for any $i=1,2,\dots,k$.
(d) This $S$ is the smallest element in $\Gamma(S)$ in terms of the partial order $\leq$. In particular, $S$ is cluster-positive.
@@
For example, by multiplying both sides of two inequatries $pqr\geq 2p^{2}$ and $pqr\geq 2q^{2}$, we have $(pqr)^{2}\geq 4p^{2}q^{2}$, and it implies $r^{2}\geq 4$ and $r\geq 2$. Similarly, we have $p,q\geq 2$.
@@
The existence of the limit was shown in [BBH11, Thm. 5.1], but we can show that this limit is $(2,2,2)$.
@@
(a) First, we show that $S_{i}$ converges to some $L(S)=(a,b,c)$ with $a,b,c\geq 2$. Let $S_{i}=(p_{i},q_{i},r_{i})$. Then, since $S_{i}$ is cluster-positive, we have $p_{i},q_{i},r_{i}\geq 2$. Moreover, these sequences $\{p_{i}\}$, $\{q_{i}\}$, and $\{r_{i}\}$ are monotonically decreasing. Thus, they converges to some real numbers $a$, $b$, and $c$, respectively. Thus, $S_{i}$ converges to $L(S)=(a,b,c)$. Moreover, since $p_{i},q_{i},r_{i}\geq 2$, we have $a,b,c\geq 2$. Next, we show $a=b=c=2$. Without loss of generality, we may assume $a\geq b\geq c$. Since monotonically decreasing sequences $p_{i}$, $q_{i}$, and $r_{i}$ converges to $a$, $b$, and $c$, respectively, for any $\varepsilon>0$, there is $N\in\mathbb{Z}_{>0}$ such that $a\leq p_{i}<a+\varepsilon$, $b\leq q_{i}<b+\varepsilon$, and $c\leq r_{i}<c+\varepsilon$ for any $i\geq N$. Since $S_{i}$ satisfies (M2), there is $i\geq N$ such that $p_{i}>q_{i}\geq r_{i}$. By Proposition 7.3 (2), we have $q_{i}r_{i}<2p_{i}$. Since $p_{i}<a+\varepsilon$ and $bc\leq q_{i}r_{i}$, we have $bc<2(a+\varepsilon)$. This inequality holds for arbtrary $\varepsilon>0$. By considering the limit $\varepsilon\to+0$, we have $bc\leq 2a$. Next, consider $S_{i+1}$. Since $p_{i}>q_{i}\geq r_{i}$ and $S_{i}\nleq S_{i+1}$, we need to choose $S_{i+1}=\gamma_{1}(S_{i})$. In order $S_{i+1}$ to satisfy (M2), we need $q_{i+1}>p_{i+1},r_{i+1}$. Thus, by doing a similar argument for $S_{i}$, we have $ac\leq 2b$. By the two inequalities $bc\leq 2a$ and $ac\leq 2b$, we have $c^{2}\leq 4$ and, by $c\geq 2$, we have $c=2$. This implies that $a=b$. Now, we assume $a=b>2=c$. Then, we have $p_{i},q_{i}>2=r_{i}$ for any $i$, and we should choose the sequence of indices as $(1,2,1,2,\dots)$ or $(2,1,2,1,\dots)$. By Proposition 6.6, we have $p_{i}=q_{i}=a$ because $S_{i}$ converges to some nonzero number $a$. However, $S_{i}=(a,a,2)$ satisfies (M1), and this contradicts the fact that $S_{i}$ satisfies (M2). Thus, we have $a=b=c=2$.
(b) Note that $C(S)=C(S_{i})=p_{i}^{2}+q_{i}^{2}+r_{i}^{2}-p_{i}q_{i}r_{i}$ does not depend on $i$. By considering the limit of $i\to\infty$, we have $C(S)=2^{2}+2^{2}+2^{2}-2\cdot 2\cdot 2=4=C(L(S))$. ∎
@@
In [BBH11], they showed it for any integer triplet $p,q,r\in\mathbb{Z}$. This result is a generalization for the real case. The following proof is similar but slightly different to the one of [BBH11, Lem. 3.3]. In particular, the proof is different when we consider $32<r<$ 3.
@@
Since $S$ satisfies (M1), by Lemma 7.5 (a) and Proposition 7.3 (1), we have $\frac{1}{2}qr\geq p\geq q\geq r\geq 2$. Now, we fix $r\geq 2$ and we see $C(S)$ as a function of $(p,q)\in\mathbb{R}^{2}$ satisfying the above inequalities. The region where $(p,q)$ satisfies these inequalities is illustrated in Figure 3. Set $f(p,q)=p^{2}+q^{2}+r^{2}-pqr$.
$p$$q$$qr=2p$$p=q$$r$$q$$(p_{1},q)$$(p,q)$$(p_{2},q)$
@@
First, we show Lemma 8.2 (1). We fix $q\geq r$ and we view $f(p,q)$ as a function of $p$. Set $p_{1}=q$ and $p_{2}=\frac{1}{2}qr$. In this case, $p$ satisfies $p_{1}\leq p\leq p_{2}$ as in Figure 3. Then, since $f(p,q)$ is a quadratic function with the positive leading coefficient $1$, we have $f(p,q)\leq\max(f(p_{1},q),f(p_{2},q))$. We can show $f(p_{1},q)\geq f(p_{2},q)$ as follows:
@@
The last inequality may be obtained by $p_{1}-p_{2}\leq 0$ and $p_{1}-qr+p_{2}=q-qr+\frac{1}{2}qr=-\frac{1}{2}q(r-2)\leq 0$. Thus, we have $f(p,q)\leq f(p_{1},q)=f(q,q)$. Moreover, since $f(q,q)=2q^{2}+r^{2}-q^{2}r=q^{2}(2-r)+r^{2}$ is a quadratic function of $q$ with the negative coefficient $02-r<$ 0 or a constant $r^{2}$, we have $f(q,q)\leq f(r,r)=C(r,r,r)$. (We used $q\geq r>0$.) Thus, we have $C(S)\leq C(r,r,r)$.
@@
9. The case (B)
Recall that there are two classes (A) and (B) of cluster-positive elements. (See Definition 7.6.) By considering Lemma 7.7 and Lemma 8.2, we have the following theorem.
Theorem 9.1.
Let $S\in\mathcal{S}^{+}$ be cluster-positive.
(a) This element $S$ satisfies (A) if and only if $C(S)<4$ or $S$ belongs to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, or $\Gamma(2,p,p)$ for some $p\geq 2$.
(b) This element $S\in\mathcal{S}^{+}$ satisfies (B) if and only if $C(S)=4$ and $S$ does not belong to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, and $\Gamma(2,p,p)$ for any $p\geq 2$.
Proof.
By Lemma 7.7, if a cluster-positive element $S$ satisfies (B), then we have $C(S)=4$. Moreover, cluster-positive elements $S\in\mathcal{S}^{+}$ satisfying $C(S)=4$ and (A) should belong to $\Gamma(p,p,2)$, $\Gamma(p,2,p)$, or $\Gamma(2,p,p)$ by Lemma 8.2 (2). Thus, the claim holds. ∎
An element $S=(\frac{29}{2},\frac{19}{4},\frac{7}{2})$ would satisfy (B). In fact, I have already checked that there is a decreasing sequence $S_{0},S_{1},\dots,S_{50}$ by using a computer. However, I cannot show that this $S$ satisfies (B). (We may check $C(S)=4$, but it would be difficult to show $S\notin\Gamma(p,p,2)\cup\Gamma(p,2,p)\cup\Gamma(2,p,p)$ for any $p\geq 2$.)
In [FT19, Rem. 4.7], another identification between (A) and (B) is shown.
@@ 10. Integer skew-symmetrizable case
Recall that $\mathcal{M}$ is defined by the set of all $M=\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)$ corresponding to a skew-symmetrizable matrix $B\in\mathrm{M}_{3}(\mathbb{R})$ as in Defnition 4.2. In this section, we consider the following subsets.
@@ 10.1. Fundamental domain
The inclusion $\widehat{\mathcal{S}}\subset\{(p,q,r)\mid p^{2},q^{2},r^{2},pqr\in\mathbb{Z}\}$ is immediately shown by the definition of $\mathrm{Sk}(M)=(p,q,r)$. Note that $pqr\in\mathbb{Z}$ follows from (10). Let $S=(p,q,r)\in\widehat{\mathcal{S}}$. We find $M\in\mathcal{M}_{\mathbb{Z}}$ with $\mathrm{Sk}(M)=S$. If at least one entry of $S$ is zero, for example $r=0$, take $M=\left(\begin{smallmatrix}p^{2}&1&0\\ 1&q^{2}&0\end{smallmatrix}\right)\in\mathcal{M}_{\mathbb{Z}}$. Assume $p,q,r\neq 0$. Since $p^{2},q^{2},r^{2}\in\mathbb{Z}$, we may express $p=l\sqrt{a}$, $q=m\sqrt{b}$, and $r=n\sqrt{c}$, where $\0l,m,n\in\mathbb{Z}\backslash\{0\}$ and $a,b,c\in\mathbb{Z}_{>0}$ do not have any square factors. Since $pqr=lmn\sqrt{abc}\in\mathbb{Z}$ and $\0lmn\in\mathbb{Z}\backslash\{0\}$, we need $\sqrt{abc}\in\mathbb{Z}$. Let $p_{k}$ be the $k$th prime number, and consider factorizations $a=\prod p_{k}^{\varepsilon_{k}^{a}}$, $b=\prod p_{k}^{\varepsilon_{k}^{b}}$, and $c=\prod p_{k}^{\varepsilon_{k}^{c}}$. Since $a$, $b$, and $c$ do not have square factors, every exponent of $p_{k}$ should be 00 or $1$. Moreover, since $\sqrt{abc}=\sqrt{\prod p_{k}^{\varepsilon_{k}^{a}+\varepsilon_{k}^{b}+\varepsilon_{k}^{c}}}$ is an integer, we need $\varepsilon_{k}^{a}+\varepsilon_{k}^{b}+\varepsilon_{k}^{c}\in 2\mathbb{Z}$ for any $k$. In particular, $(\varepsilon_{k}^{a},\varepsilon_{k}^{b},\varepsilon_{k}^{c})$ is any of $(0,0,0)$, $(1,1,0)$, $(0,1,1)$, or $(1,0,1)$. Let
@@
As Definition 7.6, every cluster-positive element $S\in\mathcal{S}$ satisfies the condition (A) or (B). However, when we consider $S\in\widehat{\mathcal{S}}$, the case (B) does not occur as follows. (This proposition is a generalization of [BBH11, Lem. 2.1 (b)] for the skew-symmetrizable case.)
@@
For this, we introduce a partial order $\leq$ on $\mathcal{M}$ as each corresponding entry satisfies this order on $\mathbb{R}$. Then, this order is compatible for the order on $\mathcal{S}$ and the map $\mathrm{Sk}$.
@@
Since $M^{\prime}\in\Gamma(M)$, there is a sequence (57) with $M_{i-1}\neq M_{i}$ and $t_{i-1}\neq t_{i}$ for any $i=1,2,\dots,k$. We show $M_{i-1}\leq M_{i}$. Consider a sequence defined by $S_{i}=\mathrm{Sk}(M_{i})$. Then, by Lemma 7.5 (c), we have $S_{i-1}\leq S_{i}$, and it implies that $M_{i-1}\leq M_{i}$ by Lemma 10.4. ∎
Theorem 10.7.
@@
=$=\left\{\left(\begin{matrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right)\in{\mathcal{M}}_{\mathbb{Z}}^{+}\mid xyz\geq 2xx^{\prime},2yy^{\prime},2zz^{\prime}\right\}$
@@
Let $M\in\mathcal{M}_{\mathbb{Z}}^{\mathrm{CP}}$. We show $\Gamma(M)\cap F\neq\emptyset$. By Proposition 10.3, there is $S=\gamma_{t}(\mathrm{Sk}(M))\in\Gamma(\mathrm{Sk}(M))$ satisfying (M1) for some finite sequence of indices. Then, since $\gamma_{t}(\mathrm{Sk}(M))=\mathrm{Sk}(\gamma_{t}(M))$, $\gamma_{t}(M)$ satisfies (M1). Thus, we have $\gamma_{t}(M)\in\Gamma(M)\cap F\neq\emptyset$. Next, we show $|\Gamma(M)\cap F|=1$. Without loss of generality, we may suppose that $M$ satisfies (M1). Let $M^{\prime}\in\Gamma(M)\cap F$. Then, by Lemma 10.6, we have $M\leq M^{\prime}$ and $M^{\prime}\leq M$, and it implies that $M=M^{\prime}$. ∎

### v2

@@ 4. Notations
Suppose that $p_{n}$ converges. This is equivalent that $f_{2k+1}$ ($k\in\mathbb{Z}_{>0}$) converges. Let $n=2k+1$. When $r=2$, it holds that $\theta=0$. Since $u_{n}(2)=U_{n}(1)=n+1$, we have $f_{n}(S)=q(n+1)-pn=(q-p)n+q$. Thus, if $f_{n}(S)$ converges, we have $q-p=qe^{0}-p=0$. By considering the maps $\gamma_{1}$ and $\gamma_{2}$, we may obtain $p_{n}=q_{n}=p=q$ for any $n\in\mathbb{Z}_{>0}$. If $r>2$, or equivalently, if $\theta>0$, it holds that $\sinh n\theta>0$ for any $n\in\mathbb{Z}_{>0}$ and we have $\frac{\sinh(n+1)\theta}{\sinh n\theta}=\frac{e^{(n+1)\theta}-e^{-(n+1)\theta}}{e^{n\theta}-e^{-n\theta}}=\frac{e^{\theta}-e^{-(2n+1)\theta}}{1-e^{-2n\theta}}\to e^{\theta}$ when $n\to\infty$. Since $u_{n}(r)=U_{n}(\cosh\theta)=\frac{\sinh(n+1)\theta}{\sinh n\theta}$, we have
@@
If $qe^{\theta}-p\neq 0$, it does not converge since $\sinh n\theta\to\infty$ when $n\to\infty$. When $p=qe^{\theta}$, we have
@@
Later, we will show $C(S)\leq 4$ for any cluster-positive element $S$. It was shown in [FT19] by using the geometrical aspects of mutations. Here, we give an alternative proof which is more closer to that in [BBH11], althoug we need some modifications for the real case.
@@
[…]
@@
First, we show that the condition $(M3)$ does not appear in $\Gamma(S)$ for any cluster-positive $S$.
@@
[…]
@@
By the above lemma, if $S$ is cluster-positive, then $\Gamma(S)$ consists of elements which satisfy either (M1) or (M2). In particular, (M1) gives some strong properties as follows.
@@
(a) We have three inequalities $p,q,r\geq 2$.
@@
which satisfies $t_{i-1}\neq t_{i}$ ($i=1,2,\dots,k$), we have $S_{i-1}\leq S_{i}$. In particular, for any $S^{\prime}\in\Gamma(S)$, there is a sequence (45) with $S_{i-1}\leq S_{i}$, and $S_{i-1}\neq S_{i}$ for any $i=1,2,\dots,k$.
(d) This $S$ is the minimum element in $\Gamma(S)$ in terms of the partial order $\leq$. In particular, $S$ is cluster-positive.
@@
For example, by multiplying both sides of two inequalities $pqr\geq 2p^{2}$ and $pqr\geq 2q^{2}$, we have $(pqr)^{2}\geq 4p^{2}q^{2}$, and it implies $r^{2}\geq 4$ and $r\geq 2$. Similarly, we have $p,q\geq 2$.
@@
The existence of the limit was shown in [BBH11, Thm. 5.1], but we can show that this limit is (2,2,2)(2,2,2)( 2 , 2 , 2 ).
@@
(a) First, we show that $S_{i}$ converges to some $L(S)=(a,b,c)$ with $a,b,c\geq 2$. Let $S_{i}=(p_{i},q_{i},r_{i})$. Then, since $S_{i}$ is cluster-positive, we have $p_{i},q_{i},r_{i}\geq 2$. Moreover, these sequences $\{p_{i}\}$, $\{q_{i}\}$, and $\{r_{i}\}$ are monotonically decreasing. Thus, they converges to some real numbers $a$, $b$, and $c$, respectively. Thus, $S_{i}$ converges to $L(S)=(a,b,c)$. Moreover, since $p_{i},q_{i},r_{i}\geq 2$, we have $a,b,c\geq 2$. Next, we show $a=b=c=2$. Without loss of generality, we may assume $a\geq b\geq c$. Since monotonically decreasing sequences $p_{i}$, $q_{i}$, and $r_{i}$ converges to $a$, $b$, and $c$, respectively, for any $\varepsilon>0$, there is $N\in\mathbb{Z}_{>0}$ such that $a\leq p_{i}<a+\varepsilon$, $b\leq q_{i}<b+\varepsilon$, and $c\leq r_{i}<c+\varepsilon$ for any $i\geq N$. Since $S_{i}$ satisfies (M2), there is $i\geq N$ such that $p_{i}>q_{i}\geq r_{i}$. By Proposition 7.3 (2), we have $q_{i}r_{i}<2p_{i}$. Since $p_{i}<a+\varepsilon$ and $bc\leq q_{i}r_{i}$, we have $bc<2(a+\varepsilon)$. This inequality holds for arbtrary $\varepsilon>0$. By considering the limit $\varepsilon\to+0$, we have $bc\leq 2a$. Next, consider $S_{i+1}$. Since $p_{i}>q_{i}\geq r_{i}$ and $S_{i}\nleq S_{i+1}$, we need to choose $S_{i+1}=\gamma_{1}(S_{i})$. In order $S_{i+1}$ to satisfy (M2), we need $q_{i+1}>p_{i+1},r_{i+1}$. Thus, by doing a similar argument for $S_{i}$, we have $ac\leq 2b$. By the two inequalities $bc\leq 2a$ and $ac\leq 2b$, we have $c^{2}\leq 4$ and, by $c\geq 2$, we have $c=2$. This implies that $a=b$. Now, we assume $a=b>2=c$. Then, we have $p_{i},q_{i}>r_{i}$ for enough large $i$, and we should choose the sequence of indices as (1,2,1,2,)(1,2,1,2$\dots$)( 1 , 2 , 1 , 2 , … ) or (2,1,2,1,)(2,1,2,1$\dots$)( 2 , 1 , 2 , 1 , … ). By Proposition 6.6, we have $p_{i}=q_{i}=a$ and $r_{i}=2$ because $S_{i}$ converges to some nonzero number $a$. However, $S_{i}=(a,a,2)$ satisfies (M1), and this contradicts the fact that $S_{i}$ satisfies (M2). Thus, we have $a=b=c=2$.
(b) Note that $C(S)=C(S_{i})=p_{i}^{2}+q_{i}^{2}+r_{i}^{2}-p_{i}q_{i}r_{i}$ does not depend on $i$. Moreover, the right hand side consists of the continuous function. Thus, by considering the limit $i\to\infty$, we have $C(S)=C(2,2,2)=4$. ∎
@@
In [BBH11], they showed it for any integer triplet $p,q,r\in\mathbb{Z}$. This result is a generalization for the real case. The following proof is similar but slightly different to the one of [BBH11, Lem. 3.3]. In particular, the proof is different when we consider $2<r<3$.
@@
[…]
@@
First, we show Lemma 8.2 (1). We fix $q\geq r$ and we view $f(p,q)$ as a function of $p$. Set $p_{1}=q$ and $p_{2}=\frac{1}{2}qr$. In this case, $p$ satisfies $p_{1}\leq p\leq p_{2}$ as in Figure 3. Then, since $f(p,q)$ is a quadratic function with the positive leading coefficient 111, we have $f(p,q)\leq\max(f(p_{1},q),f(p_{2},q))$. We can show $f(p_{1},q)\geq f(p_{2},q)$ as follows:
@@
[…]
@@ 9. Integer skew-symmetrizable case
Recall that $\mathcal{M}$ is defined by the set of all $M=\left(\begin{smallmatrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{smallmatrix}\right)$ corresponding to a skew-symmetrizable matrix $B\in\mathrm{M}_{3}(\mathbb{R})$ as in Defnition 4.1. In this section, we consider the following subsets.
@@ 9.1. Fundamental domain
The inclusion $\widehat{\mathcal{S}}\subset\{(p,q,r)\mid p^{2},q^{2},r^{2},pqr\in\mathbb{Z}\}$ is immediately shown by the definition of $\mathrm{Sk}(M)=(p,q,r)$. Note that $pqr\in\mathbb{Z}$ follows from (11). Let $S=(p,q,r)\in\widehat{\mathcal{S}}$. We find $M\in\mathcal{M}_{\mathbb{Z}}$ with $\mathrm{Sk}(M)=S$. If at least one entry of $S$ is zero, for example $r=0$, take $M=\left(\begin{smallmatrix}p^{2}&1&0\\ 1&q^{2}&0\end{smallmatrix}\right)\in\mathcal{M}_{\mathbb{Z}}$. Assume $p,q,r\neq 0$. Since $p^{2},q^{2},r^{2}\in\mathbb{Z}$, we may express $p=l\sqrt{a}$, $q=m\sqrt{b}$, and $r=n\sqrt{c}$, where $l,m,n\in\mathbb{Z}\backslash\{0\}$ and $a,b,c\in\mathbb{Z}_{>0}$ do not have any square factors. Since $pqr=lmn\sqrt{abc}\in\mathbb{Z}$ and $lmn\in\mathbb{Z}\backslash\{0\}$, we need $\sqrt{abc}\in\mathbb{Z}$. Let $p_{k}$ be the $k$th prime number, and consider factorizations $a=\prod p_{k}^{\varepsilon_{k}^{a}}$, $b=\prod p_{k}^{\varepsilon_{k}^{b}}$, and $c=\prod p_{k}^{\varepsilon_{k}^{c}}$. Since $a$, $b$, and $c$ do not have square factors, every exponent of $p_{k}$ should be 0 or 111. Moreover, since $\sqrt{abc}=\sqrt{\prod p_{k}^{\varepsilon_{k}^{a}+\varepsilon_{k}^{b}+\varepsilon_{k}^{c}}}$ is an integer, we need $\varepsilon_{k}^{a}+\varepsilon_{k}^{b}+\varepsilon_{k}^{c}\in 2\mathbb{Z}$ for any $k$. In particular, $(\varepsilon_{k}^{a},\varepsilon_{k}^{b},\varepsilon_{k}^{c})$ is any of (0,0,0)(0,0,0)( 0 , 0 , 0 ), (1,1,0)(1,1,0)( 1 , 1 , 0 ), (0,1,1)(0,1,1)( 0 , 1 , 1 ), or (1,0,1)(1,0,1)( 1 , 0 , 1 ). Let
@@
As Definition 7.6, every cluster-positive element $S\in\mathcal{S}$ satisfies the condition (A) or (B). However, when we consider $S\in\widehat{\mathcal{S}}$, the case (B) does not occur. (This proposition is a generalization of [BBH11, Lem. 2.1 (b)] for the skew-symmetrizable case.)
@@
For this, we introduce a partial order $\leq$ on $\mathcal{M}$ as each corresponding entry satisfies this order on $\mathbb{R}$. Then, this order is compatible for the map $\mathrm{Sk}$.
@@
Since $M^{\prime}\in\Gamma(M)$, there is a sequence (56) with $M_{i-1}\neq M_{i}$ and $t_{i-1}\neq t_{i}$ for any $i=1,2,\dots,k$. We show $M_{i-1}\leq M_{i}$. Consider a sequence $S_{0}=\mathrm{Sk}(M_{0}),S_{1}=\mathrm{Sk}(M_{1}),\dots,S_{k}=\mathrm{Sk}(M_{k})$. Then, by Lemma 7.5 (c), we have $S_{i-1}\leq S_{i}$, and it implies that $M_{i-1}\leq M_{i}$ by Lemma 9.4. ∎
As a conclusion, we may give a fundamental domain for the action $\Gamma$. (The existence and the uniqueness of the minimum element was shown in [Sev13, Thm. 1.2].)
Theorem 9.7 (cf. [Sev13, Thm. 1.2]).
@@
=$=\left\{\left.\left(\begin{matrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right)\in{\mathcal{M}}_{\mathbb{Z}}^{+}\ \right|\ xyz\geq 2xx^{\prime},2yy^{\prime},2zz^{\prime}\right\}$
@@
Let $M\in\mathcal{M}_{\mathbb{Z}}^{\mathrm{CP}}$. We show $\Gamma(M)\cap F\neq\emptyset$. By Proposition 9.3, there is $S=\gamma_{t}(\mathrm{Sk}(M))\in\Gamma(\mathrm{Sk}(M))$ satisfying (M1) after some transformation $\gamma_{t}$. Then, since $\gamma_{t}(\mathrm{Sk}(M))=\mathrm{Sk}(\gamma_{t}(M))$, $\gamma_{t}(M)$ satisfies (M1). Thus, we have $\gamma_{t}(M)\in\Gamma(M)\cap F\neq\emptyset$. Next, we show $|\Gamma(M)\cap F|=1$. Without loss of generality, we may suppose that $M$ satisfies (M1). Let $M^{\prime}\in\Gamma(M)\cap F$. Then, by Lemma 9.6, we have $M\leq M^{\prime}$ and $M^{\prime}\leq M$, and it implies that $M=M^{\prime}$. ∎